OpenStudy (anonymous):

Okay...um, I'm a little confused on how to approach this. So, the region of the solar system between Mars and Jupiter contains many asteroids that orbit the Sun. Consider an asteroid in a circular orbit of radius 4.2 10^11 m. Find the period of the orbit. Okay, I know I am supposed to use the equation: T^2=(4pi^2/GM)*r^3, but what exactly do I use for the M?

OpenStudy (anonymous):

Oh oops, that should be 4.2*10^11 meters.

OpenStudy (roadjester):

Use the mass being orbited. In this case the sun is at the center.

OpenStudy (anonymous):

Oh! I'll try that, thank you.

OpenStudy (anonymous):

okay, that did not turn out correct...but let me re-do the calculations. I may have done something wrong.

OpenStudy (roadjester):

make sure ur units are all consistent (that they cancel out)

OpenStudy (anonymous):

Gotcha. Ugh...I just hate really big numbers...

OpenStudy (roadjester):

are you actually writing out the numbers? use scientific notation

OpenStudy (anonymous):

I try doing both just to make sure it's as accurate as i can make it.

OpenStudy (anonymous):

Yeah ok, i'm still not getting it right

OpenStudy (roadjester):

mass of sun: 1.989 X10^30 kg G: 6.67 X10^ -11 radius of orbit: 4.2 X10^11 m

OpenStudy (anonymous):

Yep. i got all those values. I plugged them in and still i didn't get the correct answer for some reason....

OpenStudy (roadjester):

1.495 X 10^8

OpenStudy (anonymous):

ok i'm going to try this again...

OpenStudy (anonymous):

thank you for that. i just want to make sure i know how to get it.

OpenStudy (roadjester):

so was the answer i got correct?

OpenStudy (anonymous):

yes it was! thank you. i'm still stunned, how or what am i doing wrong?

OpenStudy (anonymous):

omg...i'm such a failure in life...i forgot to square root the freaking number...I apologize for being so bad in my algebra...

OpenStudy (roadjester):

\[\huge \sqrt{ \frac {4\pi^2{(4.22 X10^{11})^3}} {(6.67 X10^{-11})(1.989 X10^{30})} }\]

OpenStudy (roadjester):

.............

OpenStudy (anonymous):

Thank you. I'm really sorry for that! i didn't mean to waste your time